Elaina bicycles 4km/h faster than Robert. In the same time it takes Robert to bicycle 30 km, Elaina can bicycle 42 km. How fast does each bicyclist travel?
Answer by stanbon(75887)

Elaina bicycles 4km/h faster than Robert. In the same time it takes Robert to bicycle 30 km, Elaina can bicycle 42 km. How fast does each bicyclist travel?
-----------------
Robert DATA:
rate = x km/h ; dist = 30 km ; time = 30/x hr
-----
Elaina DATA:
rate = (x+4)km/h ; dist = 42 km ; time = 42/(x+4) hr
-----
Equation:
time = time
30/x = 42/(x+4)
30(x+4) = 42x
12x = 120
x = 10 km/h (Robert's rate)
x+4 = 14 km/h (Elaina's rate)
==================================
Cheers,
Stan H.
==================